# VideoGLUE: Video General Understanding Evaluation of Foundation Models

### v1

@@ 5.3 VideoGLUE score: Ranking FMs by their efficacies and efficiencies on video tasks
In Figure 3 we plot the averaged score achieved by each FM under each adaptation method, respectively, and compare their overall video understanding capabilities using the proposed VGS. The changes in FMs’ ranking by different adaptation methods (see the left panel in Figure 3) reinforce that the adaptation methods matter and should be considered an organic part of the evaluation of FMs. On the right panel of Figure 3, we notice that the video-native FMs overall outperform image-native FMs on our video understanding tasks, achieving averaged VGS $41.58$ vs. $39.35$ respectively. This is intuitive as video-native FMs probably have a smaller domain gap to our tasks and are more capable of temporal and motion reasoning, which are important cues for video understanding. Zooming in to the individual FMs, we find that VATT, a video-native FM, is at the first place with VGS $44.74$, followed by the image-native CLIP with VGS $41.1$. This suggests that in-domain pretraining yields overall the best adaptation capability to video tasks, and image-native FMs could also achieve competitive results on many but not all video understanding tasks.
6 Conclusion and Discussion
@@
In this report, we study three image-based and three video-based foundation models and their adaptation capability on general video understanding tasks. Experiments are conducted on three hallmark video tasks, eight diverse datasets with four distinct adaption methods. Our study shows existing image-based FMs performs well on some appearance-rich video datasets, while video-based FMs tend to achieve better on motional and temporal reasoning. Four studied adaption methods curve different landscape, revealing the critical role of considering adaption methods as an organic part of evaluating FMs. Finally, we propose one single metric VGS to represent the video task adaptation efficiency of FMs. We hope our research provides useful resources for evaluating and analyzing video foundation models, and address the current gap in foundation model evaluation within the video domain.
@@
We detail the datasets (Section A), models (Section B), and training setups (Section C) in the supplementary materials to improve this work’s reproducibility. Besides, Section D includes more experimental studies to strengthen the main text.
A Video understanding datasets
A.1 Appearance-focused action recognition
@@
Kinetics400 [28] (K400) is a large-scale, high-quality video dataset widely used as a standard video classification benchmark. It contains more than $250$k video clips with annotations of $400$ human daily actions. The actions are human focused and cover a broad range of classes including human-human interactions and human-object interactions. Although the video clips span $10$ seconds on average, many studies [45, 52] have pointed out the task could be easily solved on the Kinetics datasets by inferring from the static objects appeared or background environment — motion information is less important than the visual appearance. Hence, we categorize Kinetics400 as an appearance-focused action classification dataset.
Moments-in-Time [38] (MiT) is a large-scale video event classification dataset, with one million human annotated short video clips (around $3$ seconds each). The temporal span corresponds to the averaged duration of human working memory and is a temporal envelope holding meaningful actions between people, objects, and phenomena. Videos in MiT are annotated with 339 most used verbs in the English vocabulary.
A.2 Motion-focused action recognition
@@
Something-something v2 [20] (SSv2) is a collection of around $200$k videos of human performing pre-defined, basic actions with everyday objects. There are $174$ unique labels in total depicting atomic hand manipulations, like putting something into something, turning something upside down or covering something with something. This dataset benchmarks a model’s fine-grained understanding capability of object motions and scene changes by making the label space atomic-action-focused and background-invariant.
@@
Diving48 [34] (D48) is introduced to evaluate a model’s dynamic reasoning capability. The video clips in this dataset are obtained by segmenting online videos of major diving competitions. In total, there are around $18$k videos annotated with $48$ classes. Because of its standardization, the diving scenario is purposefully chosen to avoid the scene, object, and person biases.
A.3 Multi-label daily action classification
@@
Most of current action classification datasets involve video clips with a clean snapshot of a single action. In contrast, humans perform daily complex activities step-by-step, simultaneously, or in an interleaving manner. Towards more comprehensive human daily activity reasoning, Charades [46] is introduced. Different from web-collected datasets whose contents are more structured, Charades is collected by crowd-sourcing from hundreds of actors recording their videos in their own homes, acting out casual everyday activities. Charades brings in more diversity into the video classification task due to its close-to-daily-life setting. Its videos are $30$ seconds long on average and have multi-label annotations testing models’ understanding of complex daily activities with multiple steps. Charades provides $110$k videos with $157$ action classes for training and evaluation.
A.4 Temporal action localization
@@
Natural long videos contain scene changes and semantic shifts, while most of the existing video benchmarks formulate problems to focus on trimmed video clips. Such a gap introduces evaluation bias as clip-level benchmarks could not reflect a model’s temporal feature discriminativeness, which is of key importance to solve long-form video understanding tasks. To comprehend the study on foundation models’ video capabilities, we include the temporal action localization (TAL) task in our evaluation. The task of TAL is to predict not only the action labels but also each action instance’s temporal boundary in untrimmed videos. We adopt ActivityNet v1.3 [16] as the dataset for the TAL task, which contains $10,002$ untrimmed videos in training and $4,985$ in validation. The video length in this dataset is between $5$-$10$ minutes. In total, there are $200$ types of activities annotated.
A.5 Spatiotemporal action localization
@@
Spatiotemporal Action Localization (STAL) is a person-centric task that asks a system to localize actors and predict their atomic actions [6, 21] in a transitory duration.
@@
In AVA [21], $15$ minutes long movie clips are densely annotated at $1$Hz. In the key frames, every person is localized using a bounding box and labels corresponding to actions being performed by the actor. The label vocabulary consists of $80$ different atomic visual actions. There are $430$ different movies in total.
@@
AVA-Kinetics [31] follows the same labeling protocol as AVA, while its data source comes from the Kinetics700 [28] video pool. The dataset contains over $230$k clips annotated with the $80$ AVA action classes for each of the humans in key frames.
B Model details
B.1 Task head architectures
@@
In Figure 4, we plot the task heads used in our video classification and spatiotemporal action localization experiments, namely, the simple pooler head and multi-layer attention pooling head. For temporal localization, please refer to [57] for the task head’s detailed architecture.
@@
B.2 Image-to-video adaptation
@@
K400 SSv2
Method Early Late Early Late
CoCa [59] $72.7$ $61.4$ $41.5$ $33.3$
CLIP [41] $70.5$ $75.2$ $38.1$ $41.0$
FLAVA [47] $67.9$ $71.3$ $40.4$ $40.6$
Table 8: Ablation study on the temporal positional embedding for image-to-video adaption. We choose FLAVA [47] with the frozen feature setting in this experiment.
Temporal Positional VC (A) VC (M) VC (ML)
Embedding K400 MiT D48 SSv2 Charades
$71.3$ $29.7$ $41.6$ $30.3$ $10.7$
$71.3$ $29.7$ $45.9$ $40.6$ $12.6$
@@
We consider two choices, early-fusion and late-fusion, and ablate them in the frozen feature setting in Table 7. In both early-fusion and late-fusion, we first apply the projection layer on each frame independently to embed pixel patches into embedding tokens. We then average-pool the embedding tokens from nearby frames to reduce the sequence length to $n\times h\times w$. In the early-fusion setting, we pass all tokens together to the image backbone to extract video features. In late-fusion, we pass each set of $h\times w$ tokens independently to the image backbone. Empirically, we find that the FLAVA [47] and CLIP [41] models do better with late-fusion while CoCa [59] does better with early-fusion.
Furthermore, we ablate the importance of temporal information using the frozen-features from FLAVA [47]. In Table 8, we find that adding temporal positional embedding to the input is essential for D48 [34], SSv2 [20], and Charades [46] while not necessary for K400 [28] and MiT [38]. This supports our grouping that K400 and MiT are appearance-focused datasets.
Based on these findings, we use late-fusion for FLAVA [47] and CLIP [41] and early-fusion for CoCa [59]. We add learnable temporal positional embeddings for all the image-native FMs.
C Task-specific hyperparameters
@@
Config Kinetics400 Sth-sth v2 MiT Diving48 Charades
@@
random horizontal flip true false true true false
@@
RandAug (9, 0.5) (9, 0.5) - - -
MixUp 0.8 0.8 - - -
CutMix 1.0 1.0 - - -
@@
segment-based sample false true false false false
@@
random horizontal flip true true
@@
random color augmentation true true
@@
fps 15
per-clip length 16
@@
D More studies
D.1 Large model adaptations
Table 12: Evaluating large-scale FMs when using (a) frozen feature with a one-layer pooler head, and (b) low-rank adapter with frozen features. We report the Top-1 accuracy on K400, MiT, D48, SSv2 and MAP on Charades.
VC (A) VC (M) VC (ML)
Model Method K400 MiT D48 SSv2 Charades
InternVideo-L [55] frozen $78.6$ $33.7$ $69.6$ $67.4$ $20.9$
InternVideo-L [55] adapter $81.5$ $40.3$ $85.8$ $70.9$ $54.2$
VideoMAE-v2-B/DL [51] frozen $86.7$ $38.9$ $61.4$ $57.7$ $33.2$
VideoMAE-v2-B/DL [51] adapter $86.0$ $41.8$ $82.3$ $66.6$ $53.8$
VideoMAE-v2-g [51] frozen $59.7$ $20.7$ $42.5$ $44.2$ $12.7$
VideoMAE-v2-g [51] adapter $80.8$ $35.9$ $85.3$ $68.2$ $55.5$
VideoMAE-v2-g/FT [51] frozen $82.1$ $35.0$ $60.5$ $56.1$ $22.4$
VideoMAE-v2-g/FT [51] adapter $85.2$ $42.5$ $84.6$ $70.6$ $58.6$
For the completeness of this report and reader’s reference, in Table 12 we report experimental results under our settings with large FMs under two adaptation scenarios, namely, the frozen backbone with pooler head and the low-rank adapter. VideoMAE-v2-B/DL [51] denotes the ViT-B model distilled from ViT-g on the Kinetics710 datasets [^1] https://github.com/OpenGVLab/VideoMAEv2/blob/master/docs/MODEL_ZOO.md. VideoMAE-v2-g [51] is the model that pretrained on UnlabeledHybrid dataset, while VideoMAE-v2-g/FT [51] conducts further finetuning using supervised training on Kinetics710.
D.2 Sample-efficient transfer learning
Table 13: Benchmark FMs adaptation on video understanding tasks under sample-efficient transfer learning. This table shows Top-1 classification accuracy and the relative accuracy (shown in the bracket). Results are achieved by using frozen features with pooler head.
K400 SSv2
Method $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$
CoCa [59] $27.1(37.8\%)$ $48.9(67.0\%)$ $73.1$ $5.6(13.4\%)$ $20.9(50.4\%)$ $41.5$
CLIP [41] $36.9(46.2\%)$ $66.8(83.6\%)$ $79.0$ $8.7(19.3\%)$ $25.1(55.5\%)$ $45.3$
FLAVA [47] $14.4(20.2\%)$ $35.8(50.3\%)$ $71.3$ $7.2(17.7\%)$ $14.3(35.3\%)$ $40.6$
VideoMAE [18] $15.5(23.9\%)$ $32.0(49.2\%)$ $65.0$ $13.7(25.4\%)$ $30.3(56.2\%)$ $53.9$
InternVideo [55] $20.4(29.5\%)$ $50.2(72.4\%)$ $69.3$ $19.5(33.6\%)$ $41.1(70.7\%)$ $58.2$
VATT [1] $34.1(45.4\%)$ $63.7(84.8\%)$ $75.1$ $12.9(22.4\%)$ $37.6(65.0\%)$ $57.8$

### v3

@@ 4.3 Profiling foundation models for video understanding
On the right panel of Figure 3, we plot each FM’s VideoGLUE Score. We notice the average VGS for video-native and image-native FMs on our video understanding tasks are $40.67$ vs. $38.84$ respectively. Zooming in to the individual FMs, we find that VATT, a video-native FM, is at the first place with VGS $43.97$, followed by the image-native DINOv2 with VGS $40.17$. This suggests that in-domain pretraining yields overall the best adaptation capability to video tasks, and image-native FMs could also achieve competitive results on many but not all video understanding tasks.
5 Limitations
VideoGLUE serves as a comprehensive benchmark for studying and probing various video understanding capabilities of foundation models. The current task portfolio includes various unimodal action understanding tasks. We believe the scope of this work could be further extended as there are many other important video tasks not covered here, e.g. object or point-level tracking, long-term memory and forecasting. Moreover, our benchmark could be strengthened by adding multimodal tasks like video captioning and question answering, given the rise of general Vision Language Models (VLM). We chose three representative FM adaptation methods and used them to provide as uniform experiment protocols for different FMs as possible. However, some of our observations could be flipped with the evolution of FMs development and adaptation methods, which are an active research area. We proposed a scalar score, VideoGLUE Score (VGS), to capture the efficacy and efficiency of an FM on video understanding. However, VGS might be dominated by one or a few datasets — when it becomes a serious issue, we should probably improve the score and/or retire the datasets from future versions of VideoGLUE. Indeed, VGS is not a perfect score that covers all aspects of FMs in a comprehensive manner. For example, it does not account for an FM’s model size, model architecture, etc. We hope future research will lead to new metrics to complement VGS and a more comprehensive evaluation of FMs for visual tasks.
6 Conclusion
@@
In this report, we study four image-based and three video-based foundation models and their adaptation capability on general video understanding tasks. Experiments are conducted on three hallmark video tasks, eight diverse datasets with four distinct adaption methods. Our study shows existing image-based FMs performs well on some appearance-rich video datasets, while video-based FMs tend to achieve better on motion and temporal reasoning. Four studied adaption methods curve different landscape, revealing the critical role of considering adaption methods as an organic part of evaluating FMs. Finally, we propose one single metric VGS to represent the video task adaptation efficiency of FMs. We hope our research provides useful resources for evaluating and analyzing video foundation models, and address the current gap in foundation model evaluation within the video domain.
7 Acknowledgement
We would like to thank Xuhui Jia and Sergey Ioffe for reviewing and providing feedback on this paper. We thank Albert Shaw on their early investigation on the UNIT architecture. We also thank David Ross, Rahul Sukthankar, and Tomas Izo for their support and leadership on this project.
@@ 7 Acknowledgement
We first discuss the ethcial concerns and broader impact of this work (Section A). We detail the datasets (Section B), models (Section C), and training setups (Section D) in the supplementary materials to improve this work’s reproducibility. Besides, Section E includes more experimental studies to strengthen the main text.
Appendix A Ethical concern and broader impact
Ethical concern. We evaluate FMs on three video tasks, eight datasets in total. We select the tasks and datasets based on their popularity and representativeness. Although carefully designed, our benchmark inevitably inherited some ethical concerns from those datasets. For instance, many of the datasets are curated by crawling videos from the Internet, which do not proportionately represent the experiences of the global population and can potentially lead to biased evaluations of FMs. Moreover, the video datasets involve human daily activities, leading to privacy concerns about the human actors in the videos. How to evaluate FMs for video understanding in a fair and privacy-preserving manner could be an important direction for future research.
Broader impact. Our research reveals the need and tremendous opportunities to research video-first FMs by improving pretraining video data and methodologies. Our studies on different adaptation methods on versatile tasks confirms that both tasks and adaptation methods matter when it comes to the evaluation of FMs, shedding light on the already vibrant area of FM adaptations. Finally, we hope our research could inspire research on foundation models development and video understanding in general, along with their applications in the real world.
Appendix B Video understanding datasets
Table 7: Summary of dataset publishing year, venue and citations (as of October 17, 2024).
Task Dataset Year Venue Citation
VC Kinetics-400 (Kay et al., 2017) 2017 arXiv $4,558$
Moments in Time (Monfort et al., 2019) 2018 TPAMI $651$
Something-Something v2 (Goyal et al., 2017) 2017 ICCV $1,582$
Diving48 (Li et al., 2018) 2018 ECCV $369$
Charades (Sigurdsson et al., 2016) 2016 ECCV $1,415$
TAL ActivityNet v1.3 (Fabian Caba Heilbron & Niebles, 2015) 2015 CVPR $2,881$
STAL AVA v2.2 (Gu et al., 2018) 2018 CVPR $1,200$
AVA-Kinetics (Li et al., 2020) 2020 arXiv $151$
In Table 7 we show the publishing year, venues and citations to demonstrate the popularity and community acceptance of datasets of our choice. Below we provide dataset details.
B.1 Appearance-focused action recognition
@@
Kinetics-400 (Kay et al., 2017) (K400) is a large-scale, high-quality video dataset widely used as a standard video classification benchmark. It contains more than $250$K video clips with annotations of $400$ human daily actions. The actions are human focused and cover a broad range of classes including human-human interactions and human-object interactions. Although the video clips span $10$ seconds on average, many studies (Sevilla-Lara et al., 2021; Wang et al., 2018b) have pointed out the task could be easily solved on the Kinetics datasets by inferring from the static objects appeared or background environment — motion information is less important than the visual appearance. Hence, we categorize Kinetics400 as an appearance-focused action classification dataset.
Moments in Time (Monfort et al., 2019) (MiT) is a large-scale video event classification dataset, with one million human annotated short video clips (around $3$ seconds each). The temporal span corresponds to the averaged duration of human working memory and is a temporal envelope holding meaningful actions between people, objects, and phenomena. Videos in MiT are annotated with 339 most used verbs in the English vocabulary.
B.2 Motion-focused action recognition
@@
Something-Something v2 (Goyal et al., 2017) (SSv2) is a collection of around $200$K videos of human performing pre-defined, basic actions with everyday objects. There are $174$ unique labels in total depicting atomic hand manipulations, like putting something into something, turning something upside down or covering something with something. This dataset benchmarks a model’s fine-grained understanding capability of object motions and scene changes by making the label space atomic-action-focused and background-invariant.
@@
Diving48 (Li et al., 2018) (D48) is introduced to evaluate a model’s dynamic reasoning capability. The video clips in this dataset are obtained by segmenting online videos of major diving competitions. In total, there are around $18$K videos annotated with $48$ classes. Because of its standardization, the diving scenario is purposefully chosen to avoid the scene, object, and person biases.
B.3 Multi-label daily action classification
@@
Most of current action classification datasets involve video clips with a clean snapshot of a single action. In contrast, humans perform daily complex activities step-by-step, simultaneously, or in an interleaving manner. Towards more comprehensive human daily activity reasoning, Charades (Sigurdsson et al., 2016) is introduced. Different from web-collected datasets whose contents are more structured, Charades is collected by crowd-sourcing from hundreds of actors recording their videos in their own homes, acting out casual everyday activities. Charades brings in more diversity into the video classification task due to its close-to-daily-life setting. Its videos are $30$ seconds long on average and have multi-label annotations testing models’ understanding of complex daily activities with multiple steps. Charades provides $110$k videos with $157$ action classes for training and evaluation.
B.4 Temporal action localization
@@
Natural long videos contain scene changes and semantic shifts, while most of the existing video benchmarks formulate problems to focus on trimmed video clips. Such a gap introduces evaluation bias as clip-level benchmarks could not reflect a model’s temporal feature discriminativeness, which is of key importance to solve long-form video understanding tasks. To comprehend the study on foundation models’ video capabilities, we include the temporal action localization (TAL) task in our evaluation. The task of TAL is to predict not only the action labels but also each action instance’s temporal boundary in untrimmed videos. We adopt ActivityNet v1.3 (Fabian Caba Heilbron & Niebles, 2015) as the dataset for the TAL task, which contains $10,002$ untrimmed videos in training and $4,985$ in validation. The video length in this dataset is between $5$-$10$ minutes. In total, there are $200$ types of activities annotated.
B.5 Spatiotemporal action localization
@@
Spatiotemporal Action Localization (STAL) is a person-centric task that asks a system to localize actors and predict their atomic actions (Barker & Wright, 1955; Gu et al., 2018) in a transitory duration.
@@
In AVA (Gu et al., 2018), $15$ minutes long movie clips are densely annotated at $1$Hz. In the key frames, every person is localized using a bounding box and labels corresponding to actions being performed by the actor. The label vocabulary consists of $80$ different atomic visual actions. There are $430$ different movies in total.
@@
AVA-Kinetics (Li et al., 2020) follows the same labeling protocol as AVA, while its data source comes from the Kinetics700 (Kay et al., 2017) video pool. The dataset contains over $230$K clips annotated with the $80$ AVA action classes for each of the humans in key frames.
Appendix C Model details
C.1 Task head architectures
@@
In Figure 4, we plot the task heads used in our video classification and spatiotemporal action localization experiments, namely, the simple pooler head and multi-layer attention pooling head. For temporal localization, please refer to Xu et al. (2020) for the task head’s detailed architecture.
@@
C.2 Image-to-video adaptation
@@
Models K400 SSv2
Early Late Early Late
CLIP $70.5$ $75.2$ $38.1$ $41.0$
FLAVA $67.9$ $71.3$ $40.4$ $40.6$
CoCa $72.7$ $61.4$ $41.5$ $33.3$
Table 9: Ablation study on the temporal positional embedding for image-to-video adaption. We choose FLAVA (Singh et al., 2022) with the frozen feature setting in this experiment.
Adds temporal VC-A VC-M VC-ML
positional embedding? K400 MiT SSv2 D48 Charades
$71.3$ $29.7$ $30.3$ $41.6$ $10.7$
$71.3$ $29.7$ $40.6$ $45.9$ $12.6$
@@
We consider two choices, early-fusion and late-fusion, and ablate them in the frozen feature setting in Table 8. In both early-fusion and late-fusion, we first apply the projection layer on each frame independently to embed pixel patches into embedding tokens. We then average-pool the embedding tokens from nearby frames to reduce the sequence length to $n\times h\times w$. In the early-fusion setting, we pass all tokens together to the image backbone to extract video features. In late-fusion, we pass each set of $h\times w$ tokens independently to the image backbone. Empirically, we find that the FLAVA (Singh et al., 2022) and CLIP (Radford et al., 2021) models do better with late-fusion while CoCa (Yu et al., 2022) does better with early-fusion.
Furthermore, we ablate the importance of temporal information using the frozen-features from FLAVA (Singh et al., 2022). In Table 9, we find that adding temporal positional embedding to the input is essential for D48 (Li et al., 2018), SSv2 (Goyal et al., 2017), and Charades (Sigurdsson et al., 2016) while not necessary for K400 (Kay et al., 2017) and MiT (Monfort et al., 2019). This supports our grouping that K400 and MiT are appearance-focused datasets.
Based on these findings, we use late-fusion for FLAVA (Singh et al., 2022) and CLIP (Radford et al., 2021) and early-fusion for CoCa (Yu et al., 2022). We add learnable temporal positional embeddings for all the image-native FMs.
Appendix D Task-specific hyperparameters
@@
We performed a greedy search on the learning rate and weight decay in all our experiments while keeping most other hyperparameters (e.g., data augmentation magnitude, dropout rate, drop path rate, etc.) consistent across different models and datasets. Specifically, we start with learning rate $1$e-$4$ and weight decay $1$e-$5$ and uniformly sample learning rates and weight decay factors with a rate of $5$ and $10$, respectively, centered around the starting points. After the first round, we pick the best-identified learning rate and weight decay factor as the new starting point and conduct another round of sampling with a rate of $2$. We repeat another two to three rounds of hyperparameter search (with a rate of $2$) until the model’s performance converges. This process is a trade-off between computation costs and thoroughly examining an FM’s performance under each experiment setup. The search ranges for the learning rate and weight decay are [$4$e-$5$, $2.5$e-$3$] and [$1$e-$6$, $1$e-$4$], respectively. We found that the learning rate is the most crucial factor when adapting an FM to downstream video understanding tasks.
@@
Configurations K400 MiT SSv2 D48 Charades
General
@@
Random horizontal flip Yes Yes No Yes No
@@
RandAug (9, 0.5) - (9, 0.5) - -
MixUp 0.8 - 0.8 - -
CutMix 1.0 - 1.0 - -
@@
Segment-based sampling No No Yes No No
@@
Random horizontal flip Yes Yes
@@
Random color augmentation Yes Yes
@@
Frame rate (FPS) 15
Per-clip length (second) 16
@@
Appendix E More studies
E.1 Large model adaptations
Table 13: Evaluating large-scale FMs when using frozen feature with a one-layer pooler head. We report the Top-1 accuracy on K400, MiT, D48, SSv2 and MAP on Charades.
Models VC-A VC-M VC-ML
K400 MiT SSv2 D48 Charades
InternVideo-L $78.6$ $33.7$ $67.4$ $69.6$ $20.9$
VideoMAE-v2-B/DL $86.7$ $38.9$ $57.7$ $61.4$ $33.2$
VideoMAE-v2-g $59.7$ $20.7$ $44.2$ $42.5$ $12.7$
VideoMAE-v2-g/FT $82.1$ $35.0$ $56.1$ $60.5$ $22.4$
InternVideo-v2-g $85.0$ $43.0$ $61.6$ $53.1$ $40.9$
VideoPrism-g $86.6$ $44.7$ $67.4$ $66.1$ $61.0$
For the completeness of this report and reader’s reference, in Table 13 we report experimental results under our settings with large FMs under the frozen backbone with one pooler head setup.
VideoMAE-v2-B/DL (Wang et al., 2023) denotes the ViT-B model distilled from ViT-g on the Kinetics710 datasets [^1] https://github.com/OpenGVLab/VideoMAEv2/blob/master/docs/MODEL_ZOO.md. VideoMAE-v2-g (Wang et al., 2023) is the model that pretrained on UnlabeledHybrid dataset, while VideoMAE-v2-g/FT (Wang et al., 2023) conducts further finetuning using supervised training on Kinetics710. InternVideo-v2-g (Wang et al., 2024) and VideoPrism-g (Zhao et al., 2024) are two video foundation models with multi-stage pre-training on curated in-house web video data. For InternVideo-v2-g, we use their stage-2 checkpoint [^2] https://github.com/OpenGVLab/InternVideo/blob/main/InternVideo2/multi_modality/MODEL_ZOO.md. For videoPrism-g, we use their final checkpoint.
E.2 Sample-efficient transfer learning
Table 14: Benchmark FMs adaptation on video understanding tasks under sample-efficient transfer learning. This table shows Top-1 classification accuracy and the relative accuracy to training with $100\%$ data (shown in parentheses). Results are achieved by using frozen features with pooler head.
Models K400 SSv2
$1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$
CLIP $36.9$ $(46.2\%)$ $66.8$ $(83.6\%)$ $79.0$ $8.7$ $(19.3\%)$ $25.1$ $(55.5\%)$ $45.3$
FLAVA $14.4$ $(20.2\%)$ $35.8$ $(50.3\%)$ $71.3$ $7.2$ $(17.7\%)$ $14.3$ $(35.3\%)$ $40.6$
CoCa $27.1$ $(37.8\%)$ $48.9$ $(67.0\%)$ $73.1$ $5.6$ $(13.4\%)$ $20.9$ $(50.4\%)$ $41.5$
VATT $34.1$ $(45.4\%)$ $63.7$ $(84.8\%)$ $75.1$ $12.9$ $(22.4\%)$ $37.6$ $(65.0\%)$ $57.8$
VideoMAE $15.5$ $(23.9\%)$ $32.0$ $(49.2\%)$ $65.0$ $13.7$ $(25.4\%)$ $30.3$ $(56.2\%)$ $53.9$
InternVideo $20.4$ $(29.5\%)$ $50.2$ $(72.4\%)$ $69.3$ $19.5$ $(33.6\%)$ $41.1$ $(70.7\%)$ $58.2$